OpenStudy (anonymous):

Simplify the expression. -3(b – 7)

OpenStudy (anonymous):

-3b-21 #300

OpenStudy (ivettef365):

it should be -3b + 21

OpenStudy (anonymous):

haha listening to keef rn thanks!

OpenStudy (goldphenoix):

-3(b – 7) Distribute. -3 * b = -3b -3 * - 7= 21.

OpenStudy (goldphenoix):

So it is actually -3b + 21. :)

OpenStudy (anonymous):

minus 21?

OpenStudy (anonymous):

ok thats what i thought thank you

OpenStudy (goldphenoix):

No. When you do the distributive property, you'll end up with -3 * -7. A negative multiply by a negative is a positive. :) I hoe this help! :)

OpenStudy (ivettef365):

it is -3b +21 plus 21 (-) x (-) = +